I have to complete the square: r^2+8r+72=0
Answer by edjones(8007)

r^2+8r+72=0
r^2+8r =-72
r^2+8r+16=-56 [(8/2)^2=16 completes the square. add it to each side.
r+4=+-sqrt(-56) square root of each side.
r+4=+-2sqrt(14)i
r=-4+2sqrt(14)i, r=-4-2sqrt(14)i
.
Ed